how many chords can be produced by 6 points  in the circle  ?
a ) 12   b) 15   c ) 30  d 60 

Grade:12th pass

2 Answers

Sam
6 years ago
30 chords can be formed by 6 chords in a circle because each point can join with the other 5 pointe..therefore 6*5=30..It`s that simple..☺
Arun
6 years ago
Dear student
 
Chords using 6 points in a circle = 6C2
 = 6 ! / 4! * 2!
 = 30/2 = 15
 
 
Regards
Arun (askIITians forum expert)